# Spectral Differential Network Analysis for High-Dimensional Time Series

### v1

@@ Abstract
Spectral networks derived from multivariate time series data arise in many domains, from brain science to Earth science. Often, it is of interest to study how these networks change under different conditions. For instance, to better understand epilepsy, it would be interesting to capture the changes in the brain connectivity network as a patient experiences a seizure, using electroencephalography data. A common approach relies on estimating the networks in each condition and calculating their difference. Such estimates may behave poorly in high dimensions as the networks themselves may not be sparse in structure while their difference may be. We build upon this observation to develop an estimator of the difference in inverse spectral densities across two conditions. Using an $\ell_{1}$ penalty on the difference, consistency is established by only requiring the difference to be sparse. We illustrate the method on synthetic data experiments, on experiments with electroencephalography data, and on experiments with optogentic stimulation and micro-electrocorticography data.
@@ 1 Introduction
Spectral network analysis of multivariate time series plays a key role in fields ranging from oceanography and seismology to neuroscience (Laurindo et al., 2019; James et al., 2017; Bloch et al., 2022). Studying individual networks can provide insights into how features interact; however, it is often of interest to study how networks change across conditions or in response to an external intervention. In neuroscience, for example, many neurodegenerative disorders are associated with abnormal brain connectivity networks (Bloch et al., 2022). Spectral features are regularly used to interpret many types of neuroscientific data, from electroencephalography to magnetoencephalography data (Gnassounou et al., 2023; Richard et al., 2020; Dupré la Tour et al., 2018).
@@
Similar to correlation, coherence includes the indirect effects of other nodes in the network. Thus, coherence may not be an informative measure of the dependence between nodes. A more direct measure is the inverse spectral density. The inverse spectral density between nodes $i$ and $j$ is a rescaling of the coherence after removing the linear effects of all other nodes $\{k\neq i,j\}$ (Dahlhaus, 2000). Therefore, the inverse spectral density better resembles the effective connectivity (Friston, 2011) between brain regions, providing an initial understanding of how two regions may be causally related.
Modern data collection methods have facilitated the collection of data where the dimensionality is much greater than the number of data points. This setting is often referred to as high-dimensional ($p\gg n$). When analyzing high-dimensional time series data, regularization techniques, such as the LASSO, are needed to make the problem computationally and statistically tractable (Banerjee et al., 2008). Similarly, in the frequency domain, regularization can allow for better numerical stability and better performance (Böhm and von Sachs, 2009).
@@
In this paper, we propose Spectral D-trace Difference (SDD), a direct estimator of the difference in inverse spectral densities, which, to our knowledge, is the first of its kind. Our direct estimator only requires sparsity of the difference (Yuan et al., 2017), a more realistic assumption. For instance, despite identifying many nonzero coherence values in resting state and stimulation states, Bloch et al. (2022) found many coherence changes to be near zero. These small differences are likely due to noise in the estimation procedure, corresponding to no underlying change in coherence. Assuming sparsity of the difference, and only under a geometrically decaying time dependence condition, we show that our direct difference estimator consistently estimates the true difference. Performance of our estimator is studied in a variety of simulation settings, as well as two real data applications.
@@ 2 Method
Suppose we have mean-zero stationary data in two conditions, denoted as 1 and 2. In condition 1, there are $n_{1}$ observations of $p$ covariates, while condition 2 has $n_{2}$ observations. For example, in neuroscience applications studying how stimulation changes the brain network as in Bloch et al. (2022), conditions 1 and 2 might represent pre- and post-stimulation, respectively.
@@
The spectral density in condition $l\in\{1,2\}$ at frequency $\lambda$, $f_{l}(\lambda)$, is defined as the Fourier transform of the autocovariance matrix. Intuitively, the spectral density represents the portion of the covariance of the signal that can be attributed to the specific frequency $\lambda$ (Brockwell and Davis, 1991, pp. 331–332). In practice, many different $\lambda$ values are of interest, but for notational simplicity we assume $\lambda$ is fixed throughout and suppress the dependence on $\lambda$. Mathematically we can write our data as $\{\mathbf{x}_{l,t}\}$ for $t=1,\dots,n_{l}$ where $\mathbf{x}_{l,t}$ is a $p$-dimensional vector in condition $l\in\{1,2\}$. For each condition $l$, the autocovariance at lag $h\in\mathbb{Z}$ is denoted as $\Gamma_{l}(h)=\mathbb{E}\left(\mathbf{x}_{l,t}\mathbf{x}^{T}_{l,t+h}\right)$. If $\sum_{h=-\infty}^{\infty}\left|\Gamma_{l}(h)\right|<\infty$ then the spectral density in condition $l$ at frequency $\lambda$ exists and is defined as
@@
The spectral density, $f_{l}$, is complex-valued in general which can complicate the estimation. To transform this problem to the real space, we note that any complex matrix can be written as the sum of a real matrix and $i$ times another real matrix (for $i^{2}=-1$). Specifically we write the spectral density and its inverse as $f_{l}=A_{l}+iB_{l}$, and $f_{l}^{-1}=\tilde{A}_{l}+i\tilde{B}_{l}$ respectively where $A$, $B$, $\tilde{A}$, and $\tilde{B}$ are real matrices.
With these representations, we can now work in the real space by studying $\Sigma_{l}=\begin{bmatrix}A_{l}&-B_{l}\\ B_{l}&A_{l}\end{bmatrix}\in\mathbb{R}^{2p\times 2p}$ instead of $f_{l}$. Then, by Lemma A.1 of Fiecas et al. (2019),
@@
where $I_{2p}$ represents the $2p\times 2p$ identity matrix. It is worth noting that $\Sigma_{l}^{-1}$ is the expansion of $f_{l}^{-1}$ to the real space. Moreover, $f_{l}$ and $f_{l}^{-1}$ can easily be recovered by taking the (1,1) and (2,1) blocks of $\Sigma_{l}$ and $\Sigma_{l}^{-1}$, respectively.
@@
To justify this estimation strategy, in Appendix A.1.1 we show that this loss is convex and the population loss is minimized at the true $\Delta=\Sigma_{1}^{-1}-\Sigma_{2}^{-1}$. The inputs to this estimator are $\hat{\Sigma}_{1}$ and $\hat{\Sigma}_{2}$ which are formed by expanding $\hat{f}_{1}$ and $\hat{f}_{2}$ to the real space respectively. Our estimator is
$\hat{\Delta}=\operatorname*{arg\,min}_{\Delta}\text{\hskip 3.55658pt}\frac{1}{4}\left(\langle\hat{\Sigma}_{2}\Delta,\Delta\hat{\Sigma}_{1}\rangle+\langle\hat{\Sigma}_{1}\Delta,\Delta,\hat{\Sigma}_{2}\rangle\right)-\langle\Delta,\hat{\Sigma}_{2}-\hat{\Sigma}_{1}\rangle+\tau_{n_{1},n_{2}}\|\Delta\|_{1}$ (1)
where $\tau_{n_{1},n_{2}}$ is a penalty parameter depending on sample sizes $n_{1},n_{2}$ and $\|\Delta\|_{1}$ is the $\ell_{1}$ penalty which returns the sum of the absolute values of all entries of a matrix. We solve for $\hat{\Delta}$ using an alternating direction method of multipliers (ADMM) algorithm similar to Yuan et al. (2017).
@@ 3 Theory
In this section, we study the theoretical properties of our SDD estimator, $\hat{\Delta}$. Specifically, we establish the convergence rate of $\hat{\Delta}$ to the population quantity $\Delta$. It is easily seen in Equation 1 that our estimator $\hat{\Delta}$ relies on expanded estimates of the spectral density $\hat{\Sigma}_{1},\hat{\Sigma}_{2}$. Thus, it is natural that the rate of convergence of $\hat{\Delta}$ to $\Delta$ relies on the rates of convergence of $\hat{f}_{1}$ and $\hat{f}_{2}$ to $f_{1}$ and $f_{2}$.
@@
Suppose Assumption 2 in Appendix A.2 holds and that $\tau_{n_{1},n_{2}}\geq O(\min(n_{1},n_{2})^{-1/3})$ and $M_{n_{l}}=O(n_{l}^{2/3})$ for $n_{1}$ and $n_{2}$ large enough. Then, for any $H>0$, there exists some $C>0$ such that with probability greater than $1-Cp^{2}\min(n_{1},n_{2})^{-H}$,
$\|\hat{\Delta}-\Delta\|_{F}\leq s_{\Delta}O\left(\min(n_{1},n_{2})^{-1/3}\right),$
@@
To establish convergence rates of $\hat{f}_{1}$ and $\hat{f}_{2}$ in high-dimensions, we rely on Theorem 3.1 of Fiecas et al. (2019), which uses the functional dependency framework of Wu (2005) and Wu and Zaffaroni (2018). Specifically, our convergence rates hinge on the assumption of a geometrically decaying dependence in the data. An overview of the functional dependence framework is given in Appendix A.2 and a formal statement of the geometric decay assumption is given in Assumption 2 in Appendix A.2. The geometric decay assumption is mild and is satisfied by many processes including ARMA, ARMA-ARCH, and ARMA-GARCH processes, as well as other nonlinear autoregressive processes (Shao and Wu, 2007; Liu and Wu, 2010). A sketch of the proof for Theorem 1 is provided below while the full proof is given in Appendix A.3.
Proof sketch.
The concentration bound in Theorem 3.1 of Fiecas et al. (2019) is essential to our theoretical analysis and depends on the smoothing span $M_{n_{l}}$. We therefore begin by selecting $M_{n_{l}}$ to optimize this bound which gives $M_{n_{l}}=O(n_{l}^{2/3})$. We can then plug in this $M_{n_{l}}$ to get a version of Theorem 3.1 that only depends on constants and $n_{l}$. We let $\Lambda_{l}=\{\lambda_{j}=2\pi j/n_{l},\lfloor(n_{l}-1)/2\rfloor\leq j\leq\lfloor n_{l}/2\rfloor\}$. For each condition $l$, Theorem 3.1 of Fiecas et al. (2019) becomes, for any $H>0$ there exists a $C>0$,
$\mathbb{P}\left(\sup_{\lambda_{j}\in\Lambda_{l}}\|\hat{f}_{l}(\lambda_{j})-f_{l}(\lambda_{j})\|_{\infty}>O\big{(}n_{l}^{-1/3}\big{)}\right)\leq Cp^{2}n_{l}^{-H}\,$ (2)
where $\|A\|_{\infty}$ is the infinity norm of $A$. We proceed by using the high-dimensional M-estimator framework of Negahban et al. (2012) and similar strategies to Wang et al. (2021). In this framework, the D-trace loss in Equation 1 is our convex differentiable loss function and the $\ell_{1}$ norm is our regularization penalty. Equipped with our new concentration bound in (2), we can now prove everything needed for Corollary 1 of Negahban et al. (2012). We briefly restate some notation used below. $\|A\|_{1}$ and $\|A\|_{\infty}$ denote the $\ell_{1}$ and infinity norms of a matrix $A$ respectively while $A\otimes B$ represents the Kronecker product between $A$ and $B$. We also use $\|\mathbf{x}\|_{1}$ and $\|\mathbf{x}\|_{2}$ to denote the $\ell_{1}$ and $\ell_{2}$ norms of a vector $\mathbf{x}$.
First, we must show the restricted strong convexity (RSC) of the D-trace loss. This amounts to showing that $\mathbf{m}^{T}\left(\nabla^{2}L(\Delta,\hat{\Sigma}_{1},\hat{\Sigma}_{2})\right)\mathbf{m}>\kappa_{L}\|\mathbf{m}\|_{2}^{2}$ for all $\mathbf{m}\in\mathcal{C}(S_{\Delta})$, where $\mathcal{C}(S_{\Delta})=\left\{\mathbf{m}\in\mathbb{R}^{4p^{2}}\big{|}\|\mathbf{m}_{S_{\Delta}^{c}}\|_{1}\leq 3\|\mathbf{m}_{S_{\Delta}}\|_{1}\right\}$ for some $\kappa_{L}$. Recall that $\nabla^{2}L(\Delta,\hat{\Sigma}_{1},\hat{\Sigma}_{2})=0.5(\hat{\Sigma}_{1}\otimes\hat{\Sigma}_{2}+\hat{\Sigma_{2}}\otimes\hat{\Sigma}_{1})$. Thus, we study $\mathbf{m}^{T}(\hat{\Sigma}_{1}\otimes\hat{\Sigma}_{2})\mathbf{m}$. The same results hold for $\mathbf{m}^{T}(\hat{\Sigma}_{2}\otimes\hat{\Sigma}_{1})\mathbf{m}$. We hence get
$\mathbf{m}^{T}(\hat{\Sigma}_{1}\otimes\hat{\Sigma}_{2})\mathbf{m}\geq$ $\mathbf{m}^{T}(\Sigma_{1}\otimes\Sigma_{2})\mathbf{m}+\mathbf{m}^{T}\left(\hat{\Sigma}_{1}\otimes\hat{\Sigma}_{2}-\Sigma_{1}\otimes\Sigma_{2}\right)\mathbf{m}$
$\geq$ $\lambda_{\mathrm{min}}(\Sigma_{1})\lambda_{\mathrm{min}}(\Sigma_{2})\|\mathbf{m}\|_{2}^{2}-16s_{\Delta}\|\hat{\Sigma}_{1}\otimes\hat{\Sigma}_{2}-\Sigma_{1}\otimes\Sigma_{2}\|_{\infty}\|\mathbf{m}\|_{2}^{2},$
where $\lambda_{\mathrm{min}}(f_{l})$ is the minimum eigenvalue of $f_{l}$. For large enough $n_{l}$, $\|\hat{f}_{l}-f_{l}\|_{\infty}$ and hence $\|\hat{\Sigma}_{1}\otimes\hat{\Sigma}_{2}-\Sigma_{1}\otimes\Sigma_{2}\|_{\infty}$ are small by (2) with high probability. Thus for large enough $n_{1}$ and $n_{2}$, with high probability, we can ensure that $\mathbf{m}^{T}(\hat{\Sigma}_{1}\otimes\hat{\Sigma}_{2})\mathbf{m}\geq\frac{\lambda_{\mathrm{min}}(f_{1})\lambda_{\mathrm{min}}(f_{2})}{2}\|\mathbf{m}\|_{2}^{2}$ showing that RSC holds with $\kappa_{L}=\frac{\lambda_{\mathrm{min}}(f_{1})\lambda_{\mathrm{min}}(f_{2})}{2}$.
Next, we choose the penalty scale $\tau_{n_{1},n_{2}}$. Following Negahban et al. (2012), we require that
$\tau_{n_{1},n_{2}}$ $\geq 2\|\nabla L_{D}(\Delta,\hat{\Sigma}_{1},\hat{\Sigma}_{2})\|_{\infty}$
$=2\|0.5\left(\hat{\Sigma}_{1}\Delta\hat{\Sigma}_{2}+\hat{\Sigma}_{2}\Delta\hat{\Sigma}_{1}\right)-\left(\hat{\Sigma}_{1}-\hat{\Sigma}_{2}\right)\|_{\infty}\,.$
Let $\Gamma=0.5\left(\Sigma_{2}\otimes\Sigma_{1}+\Sigma_{1}\otimes\Sigma_{2}\right)$ and $\hat{\Gamma}=0.5\left(\hat{\Sigma}_{2}\otimes\hat{\Sigma}_{1}+\hat{\Sigma}_{1}\otimes\hat{\Sigma}_{2}\right)$. Then, we can show that
$\left\|0.5\left(\hat{\Sigma}_{1}\Delta\hat{\Sigma}_{2}+\hat{\Sigma}_{2}\Delta\hat{\Sigma}_{1}\right)-\left(\hat{\Sigma}_{1}-\hat{\Sigma}_{2}\right)\right\|_{\infty}\leq\left\|\hat{\Gamma}-\Gamma\right\|_{\infty}\left\|\Delta\right\|_{1}+\left\|\hat{\Sigma}_{1}-\Sigma_{1}\right\|_{\infty}+\left\|\hat{\Sigma}_{2}-\Sigma_{2}\right\|_{\infty}$
Similar to RSC, we can show that, for large enough $n_{1}$ and $n_{2}$, with high probability,
$\left\|0.5\left(\hat{\Sigma}_{1}\Delta\hat{\Sigma}_{2}+\hat{\Sigma}_{2}\Delta\hat{\Sigma}_{1}\right)-\left(\hat{\Sigma}_{1}-\hat{\Sigma}_{2}\right)\right\|_{\infty}\leq O(\min(n_{1},n_{2})^{-1/3}).$ (3)
Therefore, choosing $\tau_{n_{1},n_{2}}>O(\min(n_{1},n_{2})^{-1/3})$ the penalty selection requirements are satisfied with high probability. Theorem 1 then follows from Corollary 1 of Negahban et al. (2012). ∎
@@
We now discuss interesting intermediate results from the proof sketch. In equation (2), we show that, with high probability and a smoothing span of $M_{n_{l}}=O(n^{2/3})$, the deviations of the smoothed periodogram from the true spectral density in condition $l$ are $O(n_{l}^{-1/3})$. Furthermore, Equation 3 establishes that, with high probability, the maximum norm of the derivative of the D-trace loss is $O(n_{l}^{-1/3})$.
@@
Due to the dependence in the data, the convergence rate in Theorem 1 is slower than the $n^{-1/2}$ rates for i.i.d data. In fact, as we choose $M_{n}$ to minimize the deviations in Theorem 3.1 of Fiecas et al. (2019), our rate is the best we can expect to achieve using this general dependence framework when only assuming a geometrically decaying time dependence.

### v2

@@ Abstract
Spectral networks derived from multivariate time series data arise in many domains, from brain science to Earth science. Often, it is of interest to study how these networks change under different conditions. For instance, to better understand epilepsy, it would be interesting to capture the changes in the brain connectivity network as a patient experiences a seizure, using electroencephalography data. A common approach relies on estimating the networks in each condition and calculating their difference. Such estimates may behave poorly in high dimensions as the networks themselves may not be sparse in structure while their difference may be. We build upon this observation to develop an estimator of the difference in inverse spectral densities across two conditions. Using an $\ell_{1}$ penalty on the difference, consistency is established by only requiring the difference to be sparse. We illustrate the method on synthetic data experiments and on experiments with electroencephalography data.
@@ 1 INTRODUCTION
Spectral network analysis of multivariate time series plays a key role in fields ranging from oceanography and seismology to neuroscience (Laurindo et al., 2019; James et al., 2017; Bloch et al., 2022). We use the inverse spectral density matrix as our choice of network as the $i,j$ entry more directly encodes the dependence between nodes $i$ and $j$ compared to other networks in the spectral domain such as coherence Dahlhaus (2000). Studying individual networks can provide insights into how features interact; however, it is often of interest to study how networks change across conditions or in response to an external intervention (Shojaie, 2021). In neuroscience, for example, many neurodegenerative disorders are associated with abnormal brain connectivity networks (Bloch et al., 2022). Spectral features are regularly used to interpret many types of neuroscientific data, from electroencephalography to magnetoencephalography data (Gnassounou et al., 2023; Richard et al., 2020; Dupré la Tour et al., 2018).
@@
Similar to correlation, coherence includes the indirect effects of other nodes in the network. Thus, coherence may not be an informative measure of the dependence between nodes. The inverse spectral density is a more direct measure of dependence between nodes as the inverse spectral density between nodes $i$ and $j$ is a rescaling of the coherence after removing the linear effects of all other nodes $\{k\neq i,j\}$ (Dahlhaus, 2000). Therefore, the inverse spectral density better resembles the effective connectivity between brain regions (Friston, 2011), providing an initial understanding of how two regions may be causally related.
Modern data collection methods have facilitated the collection of data, where the dimensionality is much greater than the number of data points. This setting is often referred to as high-dimensional ($p\gg n$). When analyzing high-dimensional time series data, regularization techniques, such as the LASSO, are needed to make the problem computationally and statistically tractable (Banerjee et al., 2008). In the frequency domain, regularization can allow for better numerical stability and better performance (Böhm and von Sachs, 2009).
@@
In this paper, we propose Spectral D-trace Difference (SDD), a direct estimator of the difference in inverse spectral densities, which, to our knowledge, is the first method to directly target a differential network in the spectral domain. By leveraging advances in analysis of time series data, our direct estimator only requires sparsity of the difference, which is a more realistic assumption in many settings (Wang et al., 2021). For instance, despite identifying many nonzero coherence values in resting state and stimulation states, Bloch et al. (2022) found many coherence changes to be near zero. These small differences are likely due to noise in the estimation procedure, corresponding to no underlying change in coherence. Assuming sparsity of the difference, and only under a geometrically decaying time dependence condition, we show that our direct difference estimator consistently estimates the true difference. Performance of our estimator is studied in a variety of simulation settings, as well as a real data application.
@@ 2 METHOD
Suppose we have mean-zero piecewise stationary data in two conditions, denoted as 1 and 2. In condition 1, there are $n_{1}$ observations of $p$ covariates, while condition 2 has $n_{2}$ observations. It is worth noting that the assumption of piecewise stationarity only requires stationarity within each condition and not stationarity across conditions. This may be satisfied in many neuroscience applications where experimental factors affecting the brain state are known from the experimental design. For example, in Bloch et al. (2022), conditions 1 and 2 might represent pre- and post-stimulation, respectively and stimulation times are known.
@@
The spectral density in condition $l\in\{1,2\}$ at frequency $\lambda$, $f_{l}(\lambda)$, is defined as the Fourier transform of the autocovariance matrix. Intuitively, the spectral density represents the portion of the covariance of the signal that can be attributed to the specific frequency $\lambda$ (Brockwell and Davis, 1991, pp. 331–332). Mathematically we can write our data as $\{\mathbf{x}_{l,t}\}$ for $t=1,\dots,n_{l}$ where $\mathbf{x}_{l,t}$ is a $p$-dimensional vector in condition $l\in\{1,2\}$. For each condition $l$, the autocovariance at lag $h\in\mathbb{Z}$ is denoted as $\Gamma_{l}(h)=\mathbb{E}\left(\mathbf{x}_{l,t}\mathbf{x}^{T}_{l,t+h}\right)$. If $\sum_{h=-\infty}^{\infty}\left|\Gamma_{l}(h)\right|<\infty$ then the spectral density in condition $l$ at frequency $\lambda$ exists and is defined as
@@
In practice, many different $\lambda$ values are of interest, but for notational simplicity we assume $\lambda$ is fixed throughout and suppress the dependence on $\lambda$. The spectral density, $f_{l}$, is complex-valued in general which can complicate the estimation. To transform this problem to the real space, we note that any complex matrix can be written as the sum of a real matrix and $i$ times another real matrix (for $i^{2}=-1$). Specifically we write the spectral density and its inverse as $f_{l}=A_{l}+iB_{l}$, and $f_{l}^{-1}=\tilde{A}_{l}+i\tilde{B}_{l}$, respectively, where $A_{l}$, $B_{l}$, $\tilde{A}_{l}$, and $\tilde{B}_{l}$ are real matrices.
With these representations, we can now work in the real space by studying $\Sigma_{l}=\begin{bmatrix}A_{l}&-B_{l}\\ B_{l}&A_{l}\end{bmatrix}\in\mathbb{R}^{2p\times 2p}$ instead of $f_{l}$. By Lemma A.1 of Fiecas et al. (2019),
@@
where $I_{2p}$ represents the $2p\times 2p$ identity matrix. In this representation, $\Sigma_{l}^{-1}$ is the expansion of $f_{l}^{-1}$ to the real space. Moreover, $f_{l}$ and $f_{l}^{-1}$ can easily be recovered by taking the (1,1) and (2,1) blocks of $\Sigma_{l}$ and $\Sigma_{l}^{-1}$, respectively.
@@
To justify this estimation strategy, in Appendix A.1 we show that this loss is convex and the population loss is minimized at the true $\Delta=\Sigma_{1}^{-1}-\Sigma_{2}^{-1}$. The inputs to this estimator are $\hat{\Sigma}_{1}$ and $\hat{\Sigma}_{2}$ which are formed by expanding $\hat{f}_{1}$ and $\hat{f}_{2}$ to the real space respectively. Our estimator is
[…]
where $\tau_{n_{1},n_{2}}$ is a penalty parameter depending on sample sizes $n_{1},n_{2}$ and $\|\Delta\|_{1}$ is the $\ell_{1}$ penalty which returns the sum of the absolute values of all entries of a matrix. We solve for $\hat{\Delta}$ using the alternating direction method of multipliers (ADMM) algorithm from Yuan et al. (2017). Since we are inputting matrices $\hat{\Sigma}_{i}\in\mathbb{R}^{2p\times 2p}$, the computational complexity of each iteration of this algorithm is $O(8p^{3})$ while the memory requirements are $O(4p^{2})$.
@@ 3 THEORY
In this section, we study the theoretical properties of our SDD estimator, $\hat{\Delta}$. Specifically, we establish the convergence rate of $\hat{\Delta}$ to the population quantity $\Delta$. From Equation 1, our estimator $\hat{\Delta}$ relies on expanded estimates of the spectral density $\hat{\Sigma}_{1},\hat{\Sigma}_{2}$. Thus, it is natural that the rate of convergence of $\hat{\Delta}$ to $\Delta$ relies on the rates of convergence of $\hat{f}_{1}$ and $\hat{f}_{2}$ to $f_{1}$ and $f_{2}$.
@@
Suppose Assumption 1 in Appendix B holds and that $\tau_{n_{1},n_{2}}\geq O(\min(n_{1},n_{2})^{-1/3})$ and $M_{n_{l}}=O(n_{l}^{2/3})$. For $n_{1}$ and $n_{2}$ large enough and for any $H>0$, there exists some $C>0$ such that with probability greater than $1-Cp^{2}\min(n_{1},n_{2})^{-H}$,
$\|\hat{\Delta}-\Delta\|_{F}\leq O\left(s_{\Delta}\min(n_{1},n_{2})^{-1/3}\right),$
@@
The full proof of Theorem 1 is given in Appendix C. To establish convergence rates of $\hat{f}_{1}$ and $\hat{f}_{2}$ in high-dimensions, we rely in part on Theorem 3.1 of Fiecas et al. (2019), which uses the functional dependency framework of Wu (2005) and Wu and Zaffaroni (2018). Specifically, our convergence rates hinge on the assumption of a geometrically decaying dependence in the data. An overview of the functional dependence framework is given in Appendix B and a formal statement of the geometric decay assumption is given in Assumption 1 in Appendix B. The geometric decay assumption is mild and is satisfied by many processes including ARMA, ARMA-ARCH, and ARMA-GARCH processes, as well as other nonlinear autoregressive processes (Shao and Wu, 2007; Liu and Wu, 2010).
While our method uses a similar setup to Yuan et al. (2017), the analysis of our estimator is fundamentally different in two ways. First, we establish Theorem 1 only assuming sparsity of the difference and we explicitly avoid the irrepresentability assumption of Yuan et al. (2017) which has been shown not to hold in practice (Zhao and Yu, 2006). For example, Wang et al. (2021) show through simulations that for graphs of moderate or larger dimensions, this condition almost never holds. This is true even for very sparse differences and weak data dependence. Second, in contrast to the proof of Yuan et al. (2017), which assumes i.i.d. observations, our proofs also leverage advances in theoretical analysis of high-dimensional time series to handle the temporal dependence among observations over time. It is also worth noting that Theorem 1 also applies to differences in inverse spectral densities across frequencies.
@@
We now discuss interesting intermediate results from the proof. In Lemma 1, we show that, with high probability and a smoothing span of $M_{n_{l}}=O(n_{l}^{2/3})$, the deviations of the smoothed periodogram from the true spectral density in condition $l$ are $O(n_{l}^{-1/3})$. Lemma 2 shows that with high probability, the estimated Hessian of the D-trace loss converges to the true Hessian at a rate of $O(\min(n_{1},n_{2})^{1/3})$. That is, $\|\hat{\Gamma}-\Gamma\|_{\infty}=O(\min(n_{1},n_{2})^{1/3})$ where $\hat{\Gamma}=0.5\left(\hat{\Sigma}_{1}\otimes\hat{\Sigma}_{2}+\hat{\Sigma}_{2}\otimes\hat{\Sigma}_{1}\right)$ and $\Gamma$ is the same but without the hats. Lastly, Equation C establishes that, with high probability, the maximum norm of deviations of the derivative of the D-trace loss, $\left\|0.5\left(\hat{\Sigma}_{1}\Delta\hat{\Sigma}_{2}+\hat{\Sigma}_{2}\Delta\hat{\Sigma}_{1}\right)-\left(\hat{\Sigma}_{1}-\hat{\Sigma}_{2}\right)\right\|_{\infty}$, is $O(\min(n_{1},n_{2})^{-1/3})$.
@@
Due to the dependence in the data, the convergence rate in Theorem 1 is slower than the $n^{-1/2}$ rates for i.i.d data. This is because, compared with parametric models that amount to fast decaying dependence among observations, our dependence assumption is very mild and general. In fact, as we choose $M_{n}$ to minimize the deviations in Theorem 3.1 of Fiecas et al. (2019), our rate is the best we can expect to achieve using this general dependence framework when only assuming a geometrically decaying time dependence.